OpenStudy (anonymous):

The ordered pairs (1, 81), (2, 100), (3, 121), (4, 144), and (5, 169) represent a function. What is a rule that represents this function?

OpenStudy (anonymous):

Choices

OpenStudy (campbell_st):

its the 3rd choice an easy way to do this is by substitution substitute x = 1 into each equation and evaluate y = 1^8 so y = 1 y = 8^1 so y = 8 y =(1 +8)^2 so y = 9^2 or 81 y = 8(1) so y = 8 then check the 3rd option works with x = 2 if so thats the answer.

OpenStudy (anonymous):

Thank You Alot